### v3

@@ 4.7 Prediction of gradient scales along features
To probe whether LLM conceptual representations could also recover detailed knowledge about gradient scales of concepts along various features, we used human ratings spanning various categories and features. The dataset [49] includes 52 category-feature pairs, where participants were asked to rate a concept (e.g., whale) within a certain category (e.g., animal) along multiple feature dimensions (e.g., size and danger). The ratings cover nine categories including animals, cities, clothing, mythological creatures, first names, professions, sports, weather phenomena and states of the United States, with each matched with a subset of 17 features including age, arousal, cost, danger, gender, intelligence, location (indoors versus outdoors), partisanship, religiosity, size, speed, temperature, valence, (auditory) volume, wealth, weight and wetness.
For each category-feature pair, we provided the model with two demonstrations illustrating the extreme values of a target feature within a category. We then queried the model for the rating of each concept in the category to obtain the corresponding representations. For example:
the precise size rating of ants from 1 (small, little, tiny) to 5 (large, big, huge) $\Rightarrow$ 1
the precise size rating of tigers from 1 (small, little, tiny) to 5 (large, big, huge) $\Rightarrow$ ?
Similar to previous work [49], we constructed a scale vector by subtracting the representation of the minimum extreme from the maximum (e.g., $\overrightarrow{\prime\prime\textrm{size}\prime\prime_{\textrm{animal}}}=\overrightarrow{\textrm{whale}}-\overrightarrow{\textrm{ant}}$), and compared all representations to this scale to obtain their relative feature ratings. The performance of LLM-derived representations was then evaluated through the Spearman’s rank correlation between the ratings derived from them and human ratings. For each category-feature pair, we estimated a $95\%$ confidence interval based on 10,000 bootstrap samples. To control for multiple comparisons across the 52 pairs, $P$-values were corrected using the false discovery rate (FDR) method.
@@ 4.8 Word embeddings for comparison
To validate the effectiveness of LLM-derived conceptual representations in capturing human knowledge and to explore their distinct advantages over traditional static word embeddings, we compared them with a state-of-the-art 300-dimensional static word embedding trained through fastText on Common Crawl and Wikipedia [46]. Unlike LLM representations, we used cosine similarity as the similarity measure for the word embeddings, as it consistently aligned better with human behavioral data (Supplementary Information SI 1.3).
The word embeddings were compared with human behavioral data using the same method applied to LLM conceptual representations across all experiments, except for the analysis of gradient distinctions along various features. As static word embeddings do not support context-dependent computations, we followed the procedure in previous work [49] to compute a scale vector based on several antonym pairs denoting opposite values of the target feature. For instance, the opposite values for the feature “size” were represented by $\left(\overrightarrow{\textrm{large}},\overrightarrow{\textrm{big}},\overrightarrow{\textrm{huge}}\right)$ and $\left(\overrightarrow{\textrm{small}},\overrightarrow{\textrm{little}},\overrightarrow{\textrm{tiny}}\right)$, with the scale vector $\overrightarrow{\prime\prime\textrm{size}\prime\prime}$ calculated as the average of the $3\times 3=9$ pairwise vector differences between the antonyms. The word embedding for each item was then projected onto this scale vector, and the resulting projections were correlated with human ratings for evaluation.
@@ 4.9 Encoding model of neural representations in the brain
We used the fMRI dataset from the THINGS-data collection [40] to explore whether LLM conceptual representations can map onto brain activity patterns associated with visually grounded concepts. The dataset encompasses brain imaging data from three participants exposed to 8,740 representative images of 720 concepts over 12 sessions. The concepts are sampled from the total of 1,854 concepts in THINGS and are thus relevant to the behavioral data that can be predicted by LLM conceptual representations. We obtained the neural representation of each concept for each participant by averaging over its corresponding images.
To investigate the relationship between LLM-derived conceptual representations and neural representations, we trained a linear encoding model to predict voxel activations based on the LLM-derived representations. Specifically, the activation at each voxel was modeled by $y_{v,c}=\mathbf{h}_{c}\mathbf{w}+b+\epsilon$, where $y_{v,c}$ is the activation at voxel $v$ for concept $c$, $\mathbf{h}_{c}$ is the corresponding LLM representation, $\mathbf{w}$ is a vector of regression coefficients, $b$ is a constant, and $\epsilon$ represents residual error. The parameters $\mathbf{w}$ and $b$ were estimated via regularized linear regression, minimizing the least-squares error with a ridge penalty controlled by hyperparameter $\lambda$, selected through cross-validation within the training set. Model performance was assessed through twenty-fold cross-validation with non-overlapping concepts, evaluating the correlation between predicted and observed neural responses across folds. To establish statistical significance, we generated a null distribution of randomized correlation values from 10,000 test data permutations [104]. Voxel-wise $P$-values were computed by comparing predicted correlations to the null distribution and adjusted for multiple comparisons with FDR correction. We also computed noise-ceiling-normalized $R^{2}$ for each voxel by dividing the original $R^{2}$ by the estimated noise ceiling. The noise ceilings were derived based on signal and noise variance, estimated from neural activity variability to presentations of the same concept [105] (Fig. S19).
@@ 4.10 Variance partitioning between conceptual and alternative representations
To validate the effectiveness of LLM conceptual representations in explaining neural responses, we compared them against two alternative representations: (1) fastText embeddings as in Methods 4.8 and (2) 66-dimensional similarity embeddings [40] trained on 4.10 million human odd-one-out judgments of concepts in THINGS, which align well with human similarity judgments. These allowed us to examine the additional information encoded in LLM conceptual representations that aligns with human brain activity, beyond word information and similarity.
We combined each baseline representation with LLM-derived conceptual representations and analyzed the shared and unique variance each could account for [104]. To isolate unique variance, we orthogonalized the target representation and the neural responses with respect to the alternate representation, thereby removing the shared variance from both the representation and the fMRI data. The residuals of the target representation were then used to predict the fMRI residuals, and the unique variance explained by the target representation was calculated as the $R^{2}$ using twenty-fold cross-validation. For shared variance, we first concatenated both representations, used them to predict neural responses, and calculated the $R^{2}$ in the same twenty-fold cross-validation to determine the total variance explained. Shared variance was subsequently estimated by subtracting the unique contributions of each representation from this total. For our results, we focused on voxels with a noise ceiling greater than $5\%$ and reported the noise-ceiling-normalized $R^{2}$.
@@
The supplementary information contains supplementary text, Figures S8 to S19, and Tables S1 to S2.
@@
Data and Code Availability
This work makes use of previously published datasets [42, 40, 95, 47, 49, 46]. The code and additional data for reproducing all our experiments will be made publicly accessible upon publication on GitHub and Zenodo.

### v4

@@ 4.8 Prediction of gradient scales along features
We used human ratings from a dataset spanning 52 category–feature pairs [52] (https://osf.io/5r2sz/), in which participants rated a concept (e.g., whale) within a certain category (e.g., animal) along multiple feature dimensions (e.g., size and danger). The dataset covers nine categories, each paired with a subset of 17 features. For each category–feature pair, we provided the model with two demonstrations illustrating the extreme values of a target feature within a category. We then queried the model for the rating of each concept in the category to derive its representation. Similar to previous work [52], we constructed a scale vector by subtracting the representation of the minimum extreme from the maximum (e.g., $\overrightarrow{\prime\prime\textrm{size}\prime\prime_{\textrm{animal}}}=\overrightarrow{\textrm{whale}_{\textrm{size}}}-\overrightarrow{\textrm{ant}_{\textrm{size}}}$), and compared all representations to this scale to obtain their relative feature ratings. Model performance was evaluated via Spearman’s rank correlation between model-derived and human ratings. We estimated $95\%$ confidence intervals for each category–feature pair using 10,000 bootstrap samples, and corrected $P$-values for multiple comparisons across the 52 pairs using the false discovery rate (FDR) method.
@@ 4.9 Word embeddings for comparison
For predicting human concept usage, we evaluated five widely used embeddings: (1) 300-dimensional FastText word embeddings trained on Common Crawl and Wikipedia [44] (https://fasttext.cc/docs/en/crawl-vectors.html); (2) 300-dimensional DeConf sense vectors [48], which integrate word2vec embeddings with WordNet knowledge (https://pilehvar.github.io/deconf/); (3) GloVe [45]; (4) word2vec [46]; and (5) BERT embeddings [47]. Details and results for all baselines are provided in SI Appendix. In the main text, we reported FastText as a broadly covering, representative word embedding, and DeConf for its strong performance and word-sense representation. For gradient-scale analyses, we reported the best-performing embedding (GloVe), identified in the original study [52] and confirmed here, which also provided full coverage. For each embedding, the similarity measure was selected to maximize alignment with human behavioral data (SI Appendix, section SI 1.C).
Word embeddings were evaluated using the same procedures as LLM-derived conceptual representations, except for the gradient-scale analyses, where static embeddings cannot handle contextual demonstrations. Following previous work [52], scale vectors were constructed from antonym pairs denoting opposite feature values (e.g., $\overrightarrow{\textrm{big}}$ vs. $\overrightarrow{\textrm{small}}$). Each item’s embedding was projected onto the corresponding scale vector, and the resulting projections were correlated with human ratings.
@@ 4.10 Encoding model of neural representations in the brain
We used fMRI data from the THINGS-data collection [39], available at https://openneuro.org/datasets/ds004192/versions/1.0.5, to explore whether LLM-derived conceptual representations align with neural responses to visually grounded concepts. The dataset encompasses brain imaging data from three participants exposed to 8,740 images of 720 concepts over 12 sessions. These concepts, sampled from the 1,854 in THINGS, overlap with those used in our analyses of human similarity judgments. We obtained neural representations for each concept by averaging responses across its associated images. LLM representations were derived using 24 contextual demonstrations, which our analyses showed were sufficient to yield coherent and consistently structured conceptual representations.
We trained linear encoding models to predict voxel-wise activity from LLM-derived representations. Specifically, the activation at each voxel was modeled as $y_{v,c}=\mathbf{h}_{c}\mathbf{w}+b+\epsilon$, where $y_{v,c}$ is the activation at voxel $v$ for concept $c$, $\mathbf{h}_{c}$ is the corresponding LLM representation, $\mathbf{w}$ is a vector of regression coefficients, $b$ is a constant, and $\epsilon$ is residual error. The parameters $\mathbf{w}$ and $b$ were estimated via ridge regression, with the regularization parameter selected via cross-validation within the training set. Model performance was evaluated using 20-fold cross-validation with non-overlapping concepts, measuring the correlation between predicted and observed neural responses across folds. Statistical significance was assessed by comparing predicted correlations to a null distribution of correlation values generated from 10,000 random permutations [89], with FDR correction applied to voxel-wise $P$-values. We computed noise-ceiling-normalized $R^{2}$ for each voxel by dividing the original $R^{2}$ by the estimated noise ceiling. The noise ceilings were derived based on signal and noise variance, estimated from variability in neural responses to repeated presentations of the same concept [90] (SI Appendix, Fig. S28).
@@ 4.11 Variance partitioning between conceptual and alternative representations
We compared LLM-derived conceptual representations against three alternatives: (1) a 66-dimensional similarity embedding [39] trained on 4.10 million human odd-one-out judgments of concepts in THINGS (https://osf.io/f5rn6/), which aligns well with human similarity judgments; (2) static FastText embeddings [44], chosen for their full coverage of the THINGS concepts; and (3) DeConf sense vectors [48], selected for their best performance among baselines in predicting behavioral data for the THINGS concepts. These comparisons examined the explanatory power of LLM-derived representations relative to established word embeddings and human-derived similarity spaces.
[…]
@@
The supplementary information contains supplementary text, Figures S8 to S28, and Tables S1 to S4.
@@
Data, Materials, and Software Availability
This work builds on previously published datasets and materials [40, 39, 82, 49, 52, 44, 69, 48, 45, 46, 47]. All code and supplementary data necessary to reproduce the analyses are available on GitHub (https://github.com/ningyuxu/llm_concept), and archived on Zenodo (https://doi.org/10.5281/zenodo.17092983).